Exercises: (1) Verify that addition and multiplication on 2/m2 satisfy the same properties that they do in 2. . Please explain in detail, thanks!
Answer #1

10:41 5 HO + : Then Let ā, [ { Z/m2 ā= at m2, b = b + m Z for some a, bez Nowe ā + b = (at b)+ m2 + m 2 Note that mz+mZ =(a+b